# Mod $\ell$ non-vanishing of self-dual Hecke $L$-values over CM fields and applications

### v1

@@ 5.3.1. The case $\deg{{\mathfrak{p}}}=1$
${v}_{\ell}\left(L^{{\mathrm{alg}}}(1,\lambda\nu)\right)=\sum_{\begin{subarray}{c}v|{\rm{N}}_{K/F}({\mathrm{Cond}}(\lambda))\text{ inert}\\ v\nmid{\mathfrak{p}}\end{subarray}}\mu_{\ell}(\lambda_{v})$
@@
(cf. Theorem 5.7). Thus the result follows from Proposition  5.6 by taking the initial self-dual Hecke character to be $\lambda$, $\mathfrak{p}$} to be the given prime and $\chi_{0}=1$.
5.3.2. The case $\deg{{\mathfrak{p}}}>1$
Our main result is the following.
Theorem 5.9.
Let $K$ be a CM field, $F$ the maximal totally real subfield and $\Sigma$ a CM type of $K$. Let $\lambda$ be a self-dual Hecke character over $K$ of infinity type $\Sigma+\kappa(1-c)$ for $\kappa\in{\mathbb{Z}}_{\geq 0}[\Sigma]$. Let $\ell\nmid 2D_{F}$ be a prime such that $\Sigma$ is an $\ell$-ordinary CM type. If $\varepsilon(\lambda)=-1$, suppose that $\ell\neq 3$. Let ${\mathfrak{p}}\nmid\ell D_{F}{\rm N}_{K/F}({\mathrm{Cond}}(\lambda))^{-}$ be a prime of $F$ of degree at least two. Suppose that the Assumption (A) holds. Then
${v}_{\ell}\left(L^{{\mathrm{alg}}}(1,\lambda\nu)\right)=\sum_{\begin{subarray}{c}v|{\rm{N}}_{K/F}({\mathrm{Cond}}(\lambda))\text{ inert}\\ v\nmid{\mathfrak{p}}\end{subarray}}\mu_{\ell}(\lambda_{v})$
for all except finitely many $\nu\in\Xi_{\lambda,{\mathfrak{p}}}$.
Proof.
By Theorem 5.7, we have the comparison of periods
Fix an integer $a$ such that ${\bf k}_{\ell}^{\times}[p^{\infty}]=\mu_{p^{a}}$ as in (4.12). Let $\mu:O_{{\mathfrak{p}}}/{\mathfrak{p}}^{ae}\twoheadrightarrow\mu_{p^{a}}$ be a character with $\Xi_{\lambda,{\mathfrak{p}}}^{+}\cap\Xi_{{\mathfrak{p}}}(\mu)=\infty$ and $\delta_{{\mathfrak{p}},\mu}$ be as in Definition 4.26. Then by Proposition  5.6 and (5.11) we have
${v}_{\ell}\left(L^{{\mathrm{alg}}}(1,\lambda\nu)L^{{\mathrm{alg}}}(1,\lambda\nu^{-1})\right)={v}_{\ell}(\delta_{{\mathfrak{p}},\mu})+2\sum_{\text{$v|{\rm{N}}_{K/F}({\mathrm{Cond}}(\lambda))$ inert}}\mu_{\ell}(\lambda_{v})$ (5.12)
for all except finitely many $\nu\in\Xi_{\lambda,{\mathfrak{p}}}^{+}\cap\Xi_{{\mathfrak{p}}}(\mu)$. Hence, Lemma 4.27 concludes the proof. ∎
Remark 5.10.
  • $\bullet$
    The assumption (A) holds for all but finitely many $\ell$ (cf. Remark 5.5).
  • $\bullet$
    If ${\mathfrak{p}}|D_{F}$, then Theorem 5.9 still holds for all but finitely many $\ell$ (cf. Theorem 4.31).
@@ 6.1. Simultaneous mod $p$ non-vanishing of Hecke $L$-values
for $\widehat{\chi}$ the $p$-adic avatar of $\chi$ (cf. [29, Def. 1.1]), $\omega_{F}$ the Teichmüller character over $F$ and ${\mathfrak{m}}_{p}$ the maximal ideal of $\overline{{\mathbb{Z}}}_{(p)}$.
@@
often provides a lower bound for the $p$-adic valuation of $L^{{\mathrm{alg}}}(1,\chi\mu)$ for $\nu\in\Xi_{{\mathfrak{l}}}$, which is generically an equality. The following is a main result of [25, 26, 20] (cf. [20, Thm. 1.5]).
@@
Since $\chi_{2}$ is not residually self-dual, the argument in [20, §5.3.2] shows that (6.2) holds for infinitely many $\nu\in\Xi_{\chi_{1},{\mathfrak{l}}}^{+}$, concluding the proof.
@@
    If $\chi$ is not residually self-dual, then $\mu_{p}(\chi)=0$ if for each prime ${\mathfrak{q}}\parallel{\mathrm{Cond}}(\chi)$ of $F$ non-split in $K$ and coprime to $\mathfrak{l}$}, we have $p\nmid({\rm N}_{K/{\mathbb{Q}}}({\mathfrak{q}})-1)$.
@@ 6.2.1. Set-up
At precisely this stage the argument of [28] is erroneous: it is based on the mod $p$ non-vanishing of anticyclotomic Hecke $L$-values for all but finitely many $\nu$ for a given Hecke character as originally claimed in [21, 26]. Due to a subsequently found gap in Hida’s strategy towards Zariski density of CM points on self-product of the mod $p$ Hilbert modular Shimura variety and extra conditions in its correction [25, Cor. 2.12], the mod $p$ non-vanishing is only known to hold for infinitely many $\nu$ for a given Hecke character. In particular, the intersection of the set of anticyclotomic characters $\nu$ for which the mod $p$ non-vanishing holds for a given pair of Hecke characters may well be empty. Based on our mod $p$ non-vanishing of anticyclotomic Hecke $L$-values for self-dual characters, the proof of simultaneous mod $p$ non-vanishing can still be completed, as detailed in the next subsection (cf. Theorem 6.2).
@@ 6.2.2. Main results
on $U(2,1)$ as in [28, Cor. 5.15], where $\mathfrak{n}$} is an auxiliary prime to $p$ ideal of $O$ satisfying the conditions (a1)-(a3) as in [28]. It is constructed from the pullback of a Siegel Eisenstein series on $U(2,2)$, whose constant term is essentially the Katz $p$-adic $L$-function $L_{\Sigma}(\psi)$ (cf. [28, Prop. 6.6]).
@@
In the following we give a proof based on [28] and we emphasise the utility of our simultaneous mod $p$ non-vanishing of Hecke $L$-values as in Theorem 6.2. Specifically, we give an exposition [^27] We refer the reader to [28, §7] for details. of arguments of [28] which reduce the assertion to a special case of Theorem 6.2. We follow the notation of [28].
Let $\chi=\widehat{\varepsilon}\circ\Psi$ be a specialization of $\Psi$, where $\varepsilon$ is of infinity type $k\Sigma$ for $k\geq 4$ and $\widehat{\varepsilon}$ the $p$-adic avatar. Let $E^{\text{ord}}(\chi|1,{\mathfrak{n}})$ be the corresponding specialization of ${\mathcal{E}}^{\text{ord}}(\Psi|1,{\mathfrak{n}})$. Then the Fourier–Jacobi coefficients of $E^{\text{ord}}(\chi|1,{\mathfrak{n}})$ are $p$-integral holomorphic Theta functions. The space of $p$-integral Theta functions has a basis of $p$-primitive Theta functions $\Theta$ which are eigenforms for Shintani operators (cf. [28, Prop. 7.3]). For the $p$-primitivity of Fourier–Jacobi coefficients $F_{a}^{m}$, it suffices to show that the normalized inner product
@@
Let $\Theta$ be a $p$-primitive Theta function such that it is an eigenform for Shintani operators and the associated unitary self-dual Hecke character $\kappa\nu$ has epsilon factor +1+1 (cf. [28, Prop. 7.3]), where
@@
is a fixed Hecke character as in [28, Sec. 7.1] such that $\kappa|\cdot|_{{\mathbb{A}}_{K}^{\times}}^{-1/2}$ is self-dual in the sense of (1.1), $\nu$ is an auxiliary finite order anticyclotomic character on $\Gamma_{\mathfrak{l}}\simeq{\mathbb{Z}}_{\ell}$ for an auxiliary degree one primes $\mathfrak{l}$} of $F$ split in $K$.
By [28, Thm. 7.9], we have
@@
where $I(\Theta)$ is the $U(1)$-period of $\Theta$ which is $p$-integral by the main theorem of complex multiplication and arithmetic theory of line bundles on Abelian varieties (cf. [28, Sec. 7.3]), $\overline{I(\Theta)}$ is the complex conjugate of $I(\Theta)$ and $\doteq$’ denotes equality up to a $p$-unit. Moreover, it follows from $p$-integral version of the explicit Rallis inner product formula [60, 39] for the pair $(U(1),U(1))$ that
@@
(cf. [28, Thm. 7.11]). Consequently, it suffices to show that the right hand side of
@@
as in the proof of [28, Prop. 7.13 & Prop. 7.16]. ∎
@@
It may be possible to remove the hypothesis $p\nmid\Delta$ since our non-vanishing results allow $p$ and the conductor of Hecke character to have common prime factors (cf. [28, Rmk. after Thm. 2] and Remark 6.3 (iii)).
@@
In view of the $p$-primitivity of ${\mathcal{E}}(\psi)$ as in Theorem 6.5, the Eisenstein congruence approach on the unitary group $U(2,1)$ yields: the Katz $p$-adic $L$-function divides [^28] An auxiliary Deligne–Ribet $p$-adic $L$-function also appears in the argument of [28], which is innocuous due to Wiles’ proof of main conjecture for totally real number fields. the Eisenstein ideal measuring congruences between ${\mathcal{E}}(\psi)$ and the $\Lambda$-adic cusp forms on $U(2,1)$ (cf. [28, Cor. 7.20]). On the other hand, by Ribet’s lemma, the Eisenstein ideal divides the characteristic ideal of the Selmer group (cf. [28, Cor. 8.14]). ∎

### v2

@@ 5.3. Mod $\ell$ non-vanishing
${v}_{\ell}\left(L^{{\mathrm{alg}}}(1,\lambda\nu)\right)=\sum_{\begin{subarray}{c}v|{\rm{N}}_{K/F}({\mathrm{Cond}}(\lambda))\text{ inert}\end{subarray}}\mu_{\ell}(\lambda_{v})$
@@
(cf. Theorem 5.6). Thus the result follows from Proposition  5.5 by taking the initial self-dual Hecke character to be $\lambda$, $\mathfrak{p}$} to be the given prime and $\chi_{0}=1$.
@@ 6.1. Simultaneous mod $p$ non-vanishing of Hecke $L$-values
for $\widehat{\chi}$ the $p$-adic avatar of $\chi$ (cf. [28, Def. 1.1]), $\omega_{F}$ the Teichmüller character over $F$ and ${\mathfrak{m}}_{p}$ the maximal ideal of $\overline{{\mathbb{Z}}}_{p}$.
@@
often provides a lower bound for the $p$-adic valuation of $L^{{\mathrm{alg}}}(1,\chi\mu)$ for $\nu\in\Xi_{{\mathfrak{l}}}$, which is generically an equality. The following is a main result of [24, 25, 19] (cf. [19, Thm. 1.5]).
@@
Since $\chi_{2}$ is not residually self-dual, the argument in [19, §5.3.2] shows that (6.2) holds for infinitely many $\nu\in\Xi_{\chi_{1},{\mathfrak{l}}}^{+}$, concluding the proof.
@@
    If $\chi$ is not residually self-dual, then $\mu_{p}(\chi)=0$ if for each prime ${\mathfrak{q}}\parallel{\mathrm{Cond}}(\chi)$ of $K$ non-split in $K/F$ and coprime to $\mathfrak{l}$}, we have $p\nmid({\rm N}_{K/{\mathbb{Q}}}({\mathfrak{q}})-1)$.
@@ 6.2.1. Set-up
At precisely this stage the argument of [27] is erroneous: it is based on the mod $p$ non-vanishing of anticyclotomic Hecke $L$-values for all but finitely many $\nu$ for a given Hecke character as originally claimed in [20, 25]. Due to a subsequently found gap in Hida’s strategy towards Zariski density of CM points on self-product of the mod $p$ Hilbert modular Shimura variety and extra conditions in its correction [24, Cor. 2.12], the mod $p$ non-vanishing is only known to hold for infinitely many $\nu$ for a given Hecke character. In particular, the intersection of the set of anticyclotomic characters $\nu$ for which the mod $p$ non-vanishing holds for a given pair of Hecke characters may well be empty. Based on our mod $p$ non-vanishing of anticyclotomic Hecke $L$-values for self-dual characters, the proof of simultaneous mod $p$ non-vanishing can still be completed, as detailed in the next subsection (cf. Theorem 6.2).
@@ 6.2.2. Main results
on $U(2,1)$ as in [27, Cor. 5.15], where $\mathfrak{n}$} is an auxiliary prime to $p$ ideal of $O$ satisfying the conditions (a1)-(a3) as in [27]. It is constructed from the pullback of a Siegel Eisenstein series on $U(2,2)$, whose constant term is essentially the Katz $p$-adic $L$-function $L_{\Sigma}(\psi)$ (cf. [27, Prop. 6.6]).
@@
In the following we give a proof based on [27] and we emphasise the utility of our simultaneous mod $p$ non-vanishing of Hecke $L$-values as in Theorem 6.2. Specifically, we give an exposition [^27] We refer the reader to [27, §7] for details. of arguments of [27] which reduce the assertion to a special case of Theorem 6.2. We follow the notation of [27].
Let $\chi=\widehat{\varepsilon}\circ\Psi$ be a specialization of $\Psi$, where $\varepsilon$ is of infinity type $k\Sigma$ for $k\geq 4$ and $\widehat{\varepsilon}$ the $p$-adic avatar. Let $E^{\text{ord}}(\chi|1,{\mathfrak{n}})$ be the corresponding specialization of ${\mathcal{E}}^{\text{ord}}(\Psi|1,{\mathfrak{n}})$. Then the Fourier–Jacobi coefficients of $E^{\text{ord}}(\chi|1,{\mathfrak{n}})$ are $p$-integral holomorphic Theta functions. The space of $p$-integral Theta functions has a basis of $p$-primitive Theta functions $\Theta$ which are eigenforms for Shintani operators (cf. [27, Prop. 7.3]). For the $p$-primitivity of Fourier–Jacobi coefficients $F_{a}^{m}$, it suffices to show that the normalized inner product
@@
Let $\Theta$ be a $p$-primitive Theta function such that it is an eigenform for Shintani operators and the associated unitary self-dual Hecke character $\kappa\nu$ has epsilon factor +1+1 (cf. [27, Prop. 7.3]), where
@@
is a fixed Hecke character as in [27, Sec. 7.1] such that $\kappa|\cdot|_{{\mathbb{A}}_{K}^{\times}}^{-1/2}$ is self-dual in the sense of (1.1), $\nu$ is an auxiliary finite order anticyclotomic character on $\Gamma_{\mathfrak{l}}\simeq{\mathbb{Z}}_{\ell}$ for an auxiliary degree one primes $\mathfrak{l}$} of $F$ split in $K$.
In the next paragraph, we will use the notation $\doteq$’ to denote an equality up to a $p$-unit.
By [27, Thm. 7.9], we have
@@
where $I(\Theta)$ is the $U(1)$-period of $\Theta$ which is $p$-integral by the main theorem of complex multiplication and arithmetic theory of line bundles on Abelian varieties (cf. [27, Sec. 7.3]), $\overline{I(\Theta)}$ is the complex conjugate of $I(\Theta)$. Moreover, it follows from $p$-integral version of the explicit Rallis inner product formula [58, 37] for the pair $(U(1),U(1))$ that
@@
(cf. [27, Thm. 7.11]). Consequently, it suffices to show that the right hand side of
@@
as in the proof of [27, Prop. 7.13 & Prop. 7.16]. ∎
@@
It may be possible to remove the hypothesis $p\nmid\Delta$ since our non-vanishing results allow $p$ and the conductor of Hecke character to have common prime factors (cf. [27, Rmk. after Thm. 2] and Remark 6.3 (iii)).
@@
In view of the $p$-primitivity of ${\mathcal{E}}(\psi)$ as in Theorem 6.5, the Eisenstein congruence approach on the unitary group $U(2,1)$ yields: the Katz $p$-adic $L$-function divides [^28] An auxiliary Deligne–Ribet $p$-adic $L$-function also appears in the argument of [27], which is innocuous due to Wiles’ proof of main conjecture for totally real number fields. the Eisenstein ideal measuring congruences between ${\mathcal{E}}(\psi)$ and the $\Lambda$-adic cusp forms on $U(2,1)$ (cf. [27, Cor. 7.20]). On the other hand, by Ribet’s lemma, the Eisenstein ideal divides the characteristic ideal of the Selmer group (cf. [27, Cor. 8.14]). ∎